Just before Johnson Laboratories opened for business, Howard Johnson, the owner, had the following assets and liabilities. Cash $ 39,600 Laboratory Equipment 74,700 Laboratory Supplies 5,900 Loan Payable 14,200 Accounts Payable 9,650 Determine the totals that would appear in the firm’s fundamental accounting equation.
Answer #1

Calculate total that would appear in equation :

Assets = Liabilities + Stockholder's equity
Cash + Laboratory equipment + Laboratory supplies = Loan payable + Account payable
39600 + 74700 + 5900 = 14200 + 9650 + 96350

Stockholder's equity = Total assets-Total liabilities

= (39600+74700+5900)-(14200+9650)

Stockholder's equity = $96350

Assets = $120200

Liabilities = $23850